Consider the function f(x)=(2x-1)/(x-1)

If point D has a x-coordinate 2+h, what is the y-coordinate of D?

Found 4 solutions by josgarithmetic, Alan3354, math_tutor2020, ikleyn:
Answer by josgarithmetic(39618)

JUST SUBSTITUTE and calculate.

Answer by Alan3354(69443)

There's no y in the given function.

Answer by math_tutor2020(3817)


Everywhere you see an x, replace it with (2+h)

f(x) = (2x-1)/(x-1)
f(2+h) = (2(2+h)-1)/(2+h-1)
f(2+h) = (4+2h-1)/(2+h-1)
f(2+h) = (3+2h)/(1+h)

Answer: The y coordinate is (3+2h)/(1+h)

Answer by ikleyn(52790)

.
Consider the function f(x)=(2x-1)/(x-1)

If point D has a x-coordinate 2+h, what is the y-coordinate of D?
~~~~~~~~~~~~~~~~


            This my note goes to the person who created / (composed) this  " problem ".


From the given text, it is unclear (nowhere described), how point D is associated with the function f(x).


Therefore, the question is irrelevant to the given part, and vice versa,
the given part is irrelevant to the question.


Math problems are NEVER formulated in this way (!)